# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 8.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation
The first is the linear semi-gradient method.140140140The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.141141141LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.142142142An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.143143143The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.144144144The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@
8.0.2 Hu and Yang (2025): Policy Gradient for DDC Estimation
@@
The outer loop is SMM.145145145SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.146146146The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 8.1 Dynamic Oligopoly and Strategic Interaction
8.1.1 Asker, Fershtman, Jeon, and Pakes (2020): Q-Learning in Dynamic Procurement Auctions
@@
8.1.2 Hollenbeck (2019): TD Learning for Merger Analysis with Innovation
@@
The value function update for firm $i$ is149149149This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),150150150I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.151151151The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (78), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.152152152The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 8.2 Auction Equilibria and Mechanism Design
8.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).153153153A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).154154154In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@ 8.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),155155155Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
8.2.2 Ravindranath, Feng, Wang, Zaheer, Mehta, and Parkes (2024): Fitted Policy Iteration for Combinatorial Auctions
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,156156156The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.157157157DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 8.3 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)158158158DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. Moll (2025) argues that rational expectations equilibria in heterogeneous agent models are computationally intractable and proposes reinforcement learning as a more tractable alternative for modeling how agents form beliefs and make decisions; see also the textbook treatment in Zhao (2025). This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.

### 23 Mar 2026 v4

@@ 8.1.1 TD Learning for CCP Estimation
The first is the linear semi-gradient method.150150150The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.151151151LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.152152152An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.153153153The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.154154154The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@
8.1.2 Policy Gradient for DDC Estimation
@@
The outer loop is SMM.155155155SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.156156156The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 8.2 Dynamic Oligopoly and Strategic Interaction
8.2.1 Q-Learning in Dynamic Procurement Auctions
@@
8.2.2 TD Learning for Merger Analysis with Innovation
@@
The value function update for firm $i$ is159159159This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),160160160I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.161161161The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (83), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.162162162The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 8.3 Auction Equilibria and Mechanism Design
8.3.1 RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).163163163A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).164164164In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@ 8.3.1 RL for Sequential Price Mechanisms
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),165165165Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
8.3.2 Fitted Policy Iteration for Combinatorial Auctions
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,166166166The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.167167167DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 8.4 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)168168168DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. Moll (2025) argues that rational expectations equilibria in heterogeneous agent models are computationally intractable and proposes reinforcement learning as a more tractable alternative for modeling how agents form beliefs and make decisions; see also the textbook treatment in Zhao (2025). This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.